Can you please help me solve this equation: 1/2r+3s=1,for r.


Answer by jim_thompson5910(35256)

1/2r+3s=1

r+6s=2

r = 2-6s

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------